• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Boeglin

    Log entry time 16:08:58 on June 7,2010

    Entry number 325119

    keyword=Shift Summary

    8:00 continue run 4364
    8:12 start run 4364
    9:00 talked to Hall B about chaning HWP their run finished at 10am
    --------------------------------------------
    Production runs slug 29

    4364 88.4 K 147mC
    4365 70.6 K 117mC
    4366 too short
    4367 77.4 K 130.9mC
    end of slug 29
    --------------------------------------------
    10:00 switched IHWP to IN, start slug 30
    first run 4368, compton: 22960
    --------------------------------------------
    Production runs slug 30

    4368 85.3K 142.c mC
    4370 44.9K 74 mC

    --------------------------------------------
    new compton run 22961

    11:00 lost beam due to IOC reboot
    4369 junk run, no beam

    11:10 beam back start un 4370
    11:50 stopped beam asked for escorted access to Hall

    Setting up for Moeller
    12:10 Change to empty target, turn heater on
    12:20 Change helicity frequency to 30Hz
    12:45 Turn off the dithering
    13:48 Moller magnet has been refilled

    Moller measurement is still going on,